# SEA: State-Exchange Attention for High-Fidelity Physics-Based Transformers

### v1

@@ Abstract
Current approaches using sequential networks have shown promise in estimating field variables for dynamical systems, but they are often limited by high rollout errors. The unresolved issue of rollout error accumulation results in unreliable estimations as the network predicts further into the future, with each step’s error compounding and leading to an increase in inaccuracy. Here, we introduce the State-Exchange Attention (SEA) module, a novel transformer based module enabling information exchange between encoded fields through multi-head cross-attention. The cross-field multidirectional information exchange design enables all state variables in the system to exchange information with one another, capturing physical relationships and symmetries between fields. In addition, we incorporate a ViT-like architecture to generate spatially coherent mesh embeddings, further improving the model’s ability to capture spatial dependencies in the data. This enhances the model’s ability to represent complex interactions between the field variables, resulting in improved rollout error accumulation. Our results show that the Transformer model integrated with the State-Exchange Attention (SEA) module outperforms competitive baseline models, including the PbGMR-GMUS Transformer-RealNVP and GMR-GMUS Transformer, with a reduction in error of 88% and 91%, respectively, achieving state-of-the-art performance. Furthermore, we demonstrate that the SEA module alone can reduce errors by 97% for state variables that are highly dependent on other states of the system. The repository for this work is available at: https://github.com/ParsaEsmati/SEA
@@ 1 Introduction
Transformer-based architectures, in particular, are at the forefront of advancing the sequential estimation of system variables (Sun et al., 2023; Zhao et al., 2023). The output of numerical models is similar to videos, as both involve a temporal axis and a set of cells (pixels) discretizing the spatial axis. This resemblance highlights the potential of applying techniques from vision models, to enhance the accuracy and efficiency of predicting state variables in physical simulations (Lee et al., 2024; Sun et al., 2019; Arnab et al., 2021; Bertasius et al., 2021). Inspired by the spatio-temporal cross-attention mechanisms used in vision models (Lin et al., 2022; Chen et al., 2021), we propose a State-Exchange Attention (SEA) module for physics-based transformers, designed to mimic partial differential equations (PDEs) by explicitly capturing variable coupling. The effective multidirectional information exchange between fields enables the correction of some fields by others, allowing the model to learn the complex features of the physical system. This work employs a Vision Transformer (ViT) like (Dosovitskiy et al., 2020) encoder for mesh embedding. The embedded mesh at each instance is then used in the temporal State-Exchange Attention (SEA) integrated Transformer to predict the system’s future states.
@@ 4.1 Procedure
For consistent comparison with state-of-the-art models (Sun et al., 2023; Han et al., 2022), relative mean squared error is used to quantify the errors. The model was trained on an A100 GPU for approximately 2 hours for both datasets. Furthermore, a consistent Transformer architecture was adopted in both cases, utilizing 1 layer and 8 attention heads. The embedding dimension of the model for the cylinder flow case was set to 1024, in line with the literature, while a dimension of 2048 was used for the multiphase flow case to effectively capture the interface. Full details of the configurations and datasets are provided in Appendix G and F, respectively.
@@ 6 Conclusion
The presented work introduces SEA, a novel module that enables the state variables of a physical system to exchange information within the transformer architecture. Transformers integrated with SEA demonstrated state-of-the-art performance, surpassing previously established benchmarks by other transformer-based models with over 60% error reduction. This significant improvement underscores the effectiveness of SEA in capturing complex interactions within physical systems. Future work will explore the scaling complexity of SEA to handle more state variables and its application in other domains.
@@ Appendix E Theoretical error
The coupling of variables in partial differential equations (PDEs) is a common occurrence. Many physical systems are governed by sets of linear or nonlinear coupled PDEs. Examples include Maxwell’s equations (describing electromagnetism), the Ginzburg-Landau equation (used in superconductivity and phase transitions), the Einstein field equations (in general relativity), and the Navier-Stokes equations (governing fluid dynamics).
@@ Appendix H Additional results
Figure 14: Comparison of the contour maps of the velocity component u, between the predictions and ground truth in the case of $\rho=850$
@@
Figure 15: Comparison of the contour maps of the velocity component v, between the predictions and ground truth in the case of $\rho=850$
@@
Figure 16: Comparison of the contour maps of the volume fraction ($\alpha$), and velocity components between the predictions and ground truth in the case of $\rho=850$

### v2

# SEA: State-Exchange Attention for High-Fidelity Physics Based Transformers

@@ Abstract
Current approaches using sequential networks have shown promise in estimating field variables for dynamical systems, but they are often limited by high rollout errors. The unresolved issue of rollout error accumulation results in unreliable estimations as the network predicts further into the future, with each step’s error compounding and leading to an increase in inaccuracy. Here, we introduce the State-Exchange Attention (SEA) module, a novel transformer-based module enabling information exchange between encoded fields through multi-head cross-attention. The cross-field multidirectional information exchange design enables all state variables in the system to exchange information with one another, capturing physical relationships and symmetries between fields. Additionally, we introduce an efficient ViT-like mesh autoencoder to generate spatially coherent mesh embeddings for a large number of meshing cells. The SEA integrated transformer demonstrates the state-of-the-art rollout error compared to other competitive baselines. Specifically, we outperform PbGMR-GMUS Transformer-RealNVP and GMR-GMUS Transformer, with a reduction in error of 88% and 91%, respectively. Furthermore, we demonstrate that the SEA module alone can reduce errors by 97% for state variables that are highly dependent on other states of the system. The repository for this work is available at: https://github.com/ParsaEsmati/SEA
@@ 1 Introduction
The current trend towards probabilistic models, such as diffusion models for science and physics, tends to overcome the rollout errors (Han et al., 2024; Valencia et al., 2024). However, regardless of their domain (science, video, etc.), these models can produce unrealistic motions, especially for long sequence generations, and result in a temporal limitation (Weng et al., 2024). Hence, the autoregressive generation seems unavoidable for longer sequences.
Transformer-based architectures, in particular, are at the forefront of the autoregressive generations (Sun et al., 2023; Zhao et al., 2023). By drawing a parallel between the video and spatio-temporal PDE simulations, we intend to improve the unavoidable autoregressive rollout error for long sequence generation using transformers as the building block. Hence, inspired by the spatio-temporal cross-attention mechanisms used in vision models (Lin et al., 2022; Chen et al., 2021), we propose a State-Exchange Attention (SEA) module for physics-based transformers, designed to mimic PDE by explicitly capturing variable coupling. The effective multidirectional information exchange between fields enables the correction of some fields by others, allowing the model to learn the complex features of the physical system. This work employs a Vision Transformer (ViT) like (Dosovitskiy et al., 2020) encoder for mesh embedding. The embedded mesh at each instance is then used in the temporal State-Exchange Attention (SEA) integrated Transformer to predict the system’s future states.
@@ 4.1 Procedure
For consistent comparison with state-of-the-art models (Sun et al., 2023; Han et al., 2022), relative mean squared error is used to quantify the errors. The model was trained on an A100 GPU for approximately 2 hours for both datasets. Furthermore, a consistent Transformer architecture was adopted in both cases, utilizing 1 layer and 8 attention heads. The embedding dimension of the model for the cylinder flow case was set to 1024, in line with the literature (Sun et al., 2023), while a dimension of 2048 was used for the multiphase flow case to effectively capture the interface. Full details of the configurations and datasets are provided in Appendix G and F, respectively.
@@ 6 Conclusion
The presented work introduces SEA, a novel module that enables the state variables of a physical system to exchange information within the transformer architecture. Transformers integrated with SEA demonstrated state-of-the-art performance, surpassing previously established benchmarks by other transformer-based models with improved autoregressive rollout error. The improved rollout error is an indication that the SEA module is enabling the model to learn the underlying physics. In future works, we will explore the scaling complexity of SEA to handle more state variables and its application in other domains. Furthermore, the integration of SEA with probabilistic models, such as diffusion models, will be investigated.
@@ Appendix E Theoretical error
The coupling of variables in PDEs is a common occurrence. Many physical systems are governed by sets of linear or nonlinear coupled PDEs. Examples include Maxwell’s equations (describing electromagnetism), the Ginzburg-Landau equation (used in superconductivity and phase transitions), the Einstein field equations (in general relativity), and the Navier-Stokes equations (governing fluid dynamics).
@@ Appendix H Additional results
Figure 14: Comparison of the contour maps of the velocity component $u$, between the predictions and ground truth in the case of $\rho=850$.
@@
Figure 15: Comparison of the contour maps of the velocity component $v$, between the predictions and ground truth in the case of $\rho=850$.
@@
Figure 16: Comparison of the contour maps of the volume fraction ($\alpha$), and velocity components between the predictions and ground truth in the case of $\rho=850$.